40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 7 09:00:26 2022, Paco, Update, General, Scheduled power outage recovery 
    Reply  Mon Feb 7 15:32:48 2022, Yehonathan, Update, General, Y End laser controller 
    Reply  Mon Feb 7 16:53:02 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220208_001646282.jpgPXL_20220208_001657871.jpg
       Reply  Wed Feb 9 11:56:24 2022, Anchal, Update, General, Bringing back CDS Screenshot_2022-02-09_12-11-33.png
          Reply  Wed Feb 9 13:55:05 2022, Koji, Update, General, Bringing back CDS 
       Reply  Wed Feb 9 16:43:35 2022, Paco, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
          Reply  Thu Feb 10 15:41:00 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
             Reply  Thu Feb 10 17:57:48 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) PXL_20220211_021509819.jpg
                Reply  Thu Feb 10 19:03:23 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                   Reply  Thu Feb 10 19:46:37 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                      Reply  Thu Feb 10 21:10:43 2022, Koji, Update, General, Video Mux setting reset Screenshot_2022-02-10_21-11-21.pngScreenshot_2022-02-10_21-11-54.png
             Reply  Fri Feb 11 16:09:11 2022, Anchal, Update, General, Scheduled power outage recovery - Input power increased 
                Reply  Mon Feb 14 18:31:50 2022, Paco, Update, General, Scheduled power outage recovery - IMC recovery progress 
                   Reply  Tue Feb 15 19:32:50 2022, Koji, Update, General, Scheduled power outage recovery - IMC recovery progress PXL_20220216_001731377.jpgScreen_Shot_2022-02-15_at_16.18.16.pngPXL_20220216_001727465.jpgScreen_Shot_2022-02-15_at_16.22.16.pngPXL_20220216_002229572.jpg
                      Reply  Tue Feb 15 19:40:02 2022, Koji, Update, General, IMC locking 
                         Reply  Wed Feb 16 15:19:41 2022, Anchal, Update, General, Reconfigured MC reflection path for low power 
                            Reply  Wed Feb 23 15:08:57 2022, Anchal, Update, General, Removed extra beamsplitter in MC WFS path 
                               Reply  Thu Feb 24 14:32:57 2022, Anchal, Update, General, MC RFPD DCMON channel got stuck to 0 
                                  Reply  Thu Feb 24 19:26:32 2022, Anchal, Update, General, IMC Locking 
                                     Reply  Sun Feb 27 00:37:00 2022, Koji, Update, General, IMC Locking Recovery PXL_20220226_093809056.jpgPXL_20220226_093854857.jpgPXL_20220226_100859871.jpgScreenshot_2022-02-26_01-56-31.pngScreenshot_2022-02-26_01-56-47.png
                                        Reply  Sun Feb 27 01:12:46 2022, Koji, Update, General, IMC manual alignment procedure PXL_20220226_100859871.jpg
    Reply  Fri Feb 11 11:17:00 2022, Anchal, Update, General, Scheduled power outage recovery 
       Reply  Mon Feb 14 21:03:25 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220215_025325118.jpg
Message ID: 16686     Entry time: Sun Feb 27 01:12:46 2022     In reply to: 16685
Author: Koji 
Type: Update 
Category: General 
Subject: IMC manual alignment procedure 

We expect that the MC sus are susceptible to the temperature change and the alignment drifts away with time.

Here is the proper alignment procedure.

0) Assume there is no TEM00 flash or locking, but the IMC is still flashing with higher-order modes.

1) Use the CCD camera and WFS DC spots to bring the beam to the nominal position.

2) Use only MC2 and MC3 to align the cavity to have low-order modes (TEM00,01,02 etc)

3) You should be able to lock the cavity on one of these modes. Minimize the reflection (maximize the transmission) for that mode.

4) This should allow you to jump to a better lower-order mode. Continue alignment optimization only with MC2/3 until you get TEM00.

5) Optimize the TEM00 alignment only with MC2/3

6) Look at the MC end QPD. use one of the scripts in scripts/MC/moveMC2 . Note that the spot moves opposite to the name of the scripts. i.e. MC2_spot_down moves the spot up, MC2_spot_right moved the spot left, etc...
These scripts move MC1/2/3 and try to keep the good MC transmission.

7) moveMC2 scripts are not perfect. As you use them, it makes the MC alignment gradually degraded. Use MC2 and MC3 to recover good transmission.

8) If MC2 spot is satisfactory, you are done.

-------------

Step 6-8 can be done with the WFS on. This way, you can skip step 7 as the WFS servo takes care of it. But if the spot move is too fast, the servo can't keep up with the change. If so, you have to wait for the settling of the servo. Once the spot position is satisfactory, MC servo relief should be run so that the servo offset (in actuation) can be offloaded to the bias slider.

 

Attachment 1: PXL_20220226_100859871.jpg  1.844 MB  | Hide | Hide all
PXL_20220226_100859871.jpg
ELOG V3.1.3-